OpenStudy (juan337492):

One positive integer is 7 less than twice another. The sum of their squares is 433. Find the integers. Smaller integer=? larger integer=?

OpenStudy (juan337492):

(x)^2 (2x-7)^2=433

OpenStudy (jackthegreatest):

let x = a number let 2x - 7 = another number x^2 + (2x-7)^2 = 433 x^2 + 4x^2 - 24x + 49 = 433 5x^2 - 24x +49 - 433 = 0 5x^2 - 24x -384 = 0

OpenStudy (jackthegreatest):

the sum of their squares...

OpenStudy (jackthegreatest):

but what i got is not factorable... : (

OpenStudy (jackthegreatest):

@zepdrix what do you think?

OpenStudy (juan337492):

how did you get the -24x i remember you typing -28x last time before i bumb it

OpenStudy (danjs):

call the two integers X and Y. From the words you get two equations... X = 2*Y - 7 and X^2 + y^2 = 433

OpenStudy (danjs):

put X=2Y-7 into the second one \[\large (2y-7)^2 + y^2 = 433\] you can solve that for Y, expand out the first quantity to begin, \[\large (4y^2-14y -14y +49) + y^2 = 433\]

OpenStudy (danjs):

combining those to a quadratic ax^2 + bx + c = 0 \[\large 5y^2-28y-384 = 0\]

OpenStudy (danjs):

if you cant factor that, you can always use the quadratic formula, for ax^2 + bx + c = 0 \[\large x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (danjs):

\[\large y=\frac{ 28 \pm \sqrt{28^2-4*5*-384} }{ 2*5 }\] \[\large y=12~~~~~or~~~~~y=-32/5\]

OpenStudy (danjs):

It says both integers are positive, so one of them is y=12 the other is X = 2Y-7 X = 2*12 - 7 = 17 X=17 and Y=12 test and see if both conditions hold for those values

OpenStudy (juan337492):

how did you get y=12